The inductance of a solenoid with 450 turns and a length of 24 cm is 7.3 mH. (a) What is the cross-sectional area of the solenoid? (b) What is the induced emf in the solenoid if its current drops from 3.2 A to 0 in 55 ms?
Answer #1

solution-

a) we know the inductance of the solenoid

and

where A=area

N=number of turns =450

l=length=24cm=.24m

L=inductance=7.3*10^-3H

therefore

b) induced emf E=LdI/dt

where dI/dt= change in current /time interwal=3.2A-0A/55*10^-3s=58.18A/s

and E=(7.3*10^-3)(58.18A/s)=0.425V